Mackenzie invests $800 in one account and $2400 in an account paying 3 % higher interest. At the end of one year she had earned $264 in interest. At what rates did she invest?
Answer by stanbon(75887)

Mackenzie invests $800 in one account and $2400 in an account paying 3 % higher interest. At the end of one year she had earned $264 in interest. At what rates did she invest?
---------------
interest + interest = interest
800(x/100) + 2400(x+3)/100 = 264
---------------------------
800x + 2400x + 7200 = 26400
-------------------------------
3200x = 19200
-------
x = 6%
(interest on the $800)
---------
x+3 = 9% (interest on the $2400)
-----------
Cheers,
Stan H.
-------------